OpenStudy (anonymous):

find if the sequence converges or diverges {(-1^)n (2n^2-n)/(n^3+n+1)}

OpenStudy (dumbcow):

http://www.wolframalpha.com/input/?i=%28-1%29^n+*%282n^2-n%29%2F%28n^3+%2Bn%2B1%29 nope i was wrong, it converges to 0

OpenStudy (dumbcow):

if you multiply the sequence by \[\large \frac{\frac{1}{n^{3}}}{\frac{1}{n^{3}}}\] then take limit as n->inf

OpenStudy (anonymous):

thx